betty is twice as old as cathy was two years ago. the difference of their present age is 7. find the age of betty now.
Answer by tommyt3rd(5050)

1) B=2(C-2)=2C-4
2) B-C=7

B=C+7

so that
C+7=2C-4
7+4=2C-C
C=11
thus

B=18